# GazeHTA: End-to-end Gaze Target Detection with Head-Target Association

### v2

@@ Abstract
We propose an end-to-end approach for gaze target detection: predicting a head-target connection between individuals and the target image regions they are looking at. Most of the existing methods use independent components such as off-the-shelf head detectors or have problems in establishing associations between heads and gaze targets. In contrast, we investigate an end-to-end multi-person Gaze target detection framework with Heads and Targets Association (GazeHTA), which predicts multiple head-target instances based solely on input scene image. GazeHTA addresses challenges in gaze target detection by (1) leveraging a pre-trained diffusion model to extract scene features for rich semantic understanding, (2) re-injecting a head feature to enhance the head priors for improved head understanding, and (3) learning a connection map as the explicit visual associations between heads and gaze targets. Our extensive experimental results demonstrate that GazeHTA outperforms state-of-the-art gaze target detection methods and two adapted diffusion-based baselines on two standard datasets.
1 Introduction
Human attention estimation using visual observations is critical for real-world applications such as human-robot interaction [1, 30, 32], social activity recognition [12], mental health diagnosis [3], and customer behavior analysis [38]. The common solution for human attention estimation is to infer the 3D direction of gaze from human eyes [46, 45] or face [46, 7]. However, many downstream applications require the specific region or object that a person is looking at, i.e., the gaze target, rather than the gaze direction.
@@ 1 Introduction
Gaze target detection aims to directly associate individuals and their gaze targets within a single image [25] or across multiple video frames [26], offering an end-to-end solution for human attention estimation. Most of the gaze target detection approaches employ two-stream architectures, where one stream focuses on scene feature extraction while the other one learns head features [8, 16, 9, 2]. These approaches face challenges such as the lack of direct association between heads and gaze targets, and the dependency on off-the-shelf head detectors as demonstrated in recent studies [41, 40]. Moreover, the majority of previous methods are restricted to processing one head at a time, requiring repeated processing to identify gaze targets for all individuals in the scene when multiple people are present. The work of Tu et al. [41] addresses the challenges in the two-stream approaches with an end-to-end multi-person gaze target detection method, to directly predict multiple pairs of head-target instances. Nevertheless, this approach does not establish strong connections between heads and gaze targets, which restricts the final detection accuracy. The work of Tonini et al. [40] improves the association between heads and gaze targets, and achieves state-of-the-art performance. It is, however, not fully end-to-end due to its dependency on preliminary object detection. Pre-training the object detector also limits the object categories considered for gaze targets. This constraint may exclude diverse object classes or regions within the environment that are beyond the scope of the pre-trained object detector.
@@
Thus, there are two major challenges in gaze target detection: 1) detecting the heads and the gaze targets; and 2) associating heads and gaze targets to predict which head is looking at which target. Addressing the first challenge involves identifying the relevant heads, objects or other regions as potential gaze targets. This can be addressed using powerful feature extractors, such as diffusion-based models, which have demonstrated remarkable performance in various visual perception tasks [6, 14, 27, 29, 47]. However, effectively integrating these pre-trained large foundation models into gaze target detection remains an open question. The second challenge entails linking heads and gaze targets, which requires gaze target localization while considering the presence of individual heads. Despite efforts to employ gaze estimation results as constraints on the gaze target locations [16, 13, 40, 35], most existing methods have not fully addressed the association between heads and gaze targets.
@@
In this paper, we propose GazeHTA: a novel end-to-end framework for multi-person gaze target detection from a single scene image. We exploit features from a pre-trained diffusion model, while emphasizing head and facial regions with a novel re-injection method. We then associate the heads and gaze targets explicitly by a connection map in the image. These connection maps serve as links between the heads and gaze targets in the image, offering additional supervision during model training. An overview is shown in Fig. 1. Our experiments demonstrate that GazeHTA achieves state-of-the-art performance while not being restricted to the diffusion model backbone. In summary, the main contributions of this paper include:
@@
Our extensive experiments show the proposed GazeHTA achieves state-of-the-art performances on standard datasets.
2 Related Works
2.1 Gaze Target Detection
@@
Recently, end-to-end approaches have emerged to simultaneously detect multiple heads and their associated gaze targets [41, 40]. The HGTTR is a transformer-based network that predicts several head-target instances comprising head locations, out-of-frame flags, and gaze heatmaps [41]. Unfortunately, its performance suffers from brittle associations between heads and gaze targets. As a result, the latest advancements in multi-person gaze target detection have led to the development of object-aware frameworks, which leverage the concept of objects to establish associations between heads and gaze targets [40, 42]. The object-aware frameworks initially predict objects and subsequently associate head objects with all detected objects for gaze target detection. However, these methods revert to a two-stage approach because of their reliance on detected objects for head-object association. Moreover, the scope of gaze targets is limited to a subset of classes for which the pre-trained object detector was trained, raising concerns about the robustness and applicability of these methods in real-world scenarios. In contrast to previous approaches, GazeHTA offers an end-to-end solution with implicit target feature extraction, enhanced head priors, and finally, an explicit association between heads and gaze targets through connection maps.
2.2 Diffusion Models
Originally from diffusion probabilistic models [34], diffusion models are generative models designed to learn the distribution of data for generating new samples resembling the training data. This process involves iteratively applying a series of transformations from a noise image to a realistic image [27, 23, 28]. A significant milestone in text-to-image generation has been achieved by Stable Diffusion [27] and DALL-E [24]. In particular, Stable Diffusion is released with open-source code and operates within the latent space of powerful pre-trained autoencoders, enabling training on limited computational resources while preserving quality and flexibility.
@@ 2.2 Diffusion Models
With the remarkable success of diffusion models in image generation, recent studies showed their potential across various downstream visual perception tasks [37, 10]. It is straightforward to integrate a pre-trained diffusion model within diffusion-based pipelines, where target outputs are generated by iteratively removing noise from input noise sources [6, 14, 31, 20]. This strategy can be extended to gaze target detection, where the scene image serves as the condition for predicting head and gaze heatmaps through multi-step denoising. However, text-to-image diffusion models are primarily designed to generate high-quality images from noise, while our gaze target detection task uses connection map prediction. Therefore, directly applying the denoising process may not be efficient for gaze target detection.
@@
Alternatively, pre-trained diffusion models can be used as feature extraction backbones, without engaging in the progressive denoising process [47, 5]. Despite its simplicity, employing a pre-trained diffusion model as a feature extractor is effective for various visual perception tasks, including image segmentation and depth estimation [47, 43]. Considering the output format of gaze target detection is a heatmap which resembles that of image segmentation and depth estimation, namely the spatial maps or heatmaps, they may potentially benefit from the same model architecture. Following the same strategy, GazeHTA employs a pre-trained diffusion model as the backbone to extract both high-level and low-level semantic features for object understanding and head localization, which are crucial for accurate gaze target detection.
3 GazeHTA
@@
The pipeline of GazeHTA is shown in Fig. 2. Taking only a scene image as the input, GazeHTA extracts multi-scale scene features for out-of-frame binary predictions and head-target proposals. For the head-target proposal, a head feature is learned and re-injected back to enhance the head priors. Subsequently, heads, gaze targets, and connection maps are predicted in the form of heatmaps. Within each head-target instance, GazeHTA establishes a visual association between the head and its corresponding gaze target by a connection map.
3.1 Scene Feature Extraction
@@
In GazeHTA, the backbone to extract scene features is a pre-trained denoising U-Net from Stable Diffusion [27]. The use of Stable Diffusion, a large-scale text-to-image generator, is due to its capability to encompass features with both low-level and high-level semantic knowledge [22] acquired from large-scale real-world data. By taking the scene image as the only input, GazeHTA encodes it with a pre-trained image encoder, followed by a denoising U-Net from the pre-trained diffusion model. We extract intermediate features from four different levels in the U-Net with a similar operation to the VPD method [47]. These multi-scale scene features are then encoded into a compact representation, denoted as . is utilized for both out-of-frame predictions and head-target proposals. We expect the implicit understanding of the scene features can facilitate the identification of potential heads and gaze targets.
Note that, unlike VPD [47], we eliminate the entire text prompt module by removing the cross-attention layers in the denoising U-Net, as it is not necessary nor effective for the gaze target detection task.
3.2 Head Feature Re-injection
@@
Accurately locating the human head position is essential for successful gaze target detection. To improve head detection, we introduce supervision on a head detection map, which aims to detect all instances of heads within the scene image. To this end, the encoded feature is passed through a convolution-based decoder that gradually expands the spatial dimension of into a decoded feature, denoted as . is then passed through convolutional layers to generate an intermediate head detection map for head prior enhancement. To supervise the training on , we apply the detection results from the RetinaFace [11] as pseudo labels since the heads are not fully annotated in the training data. Although provides insights into potential head positions, we only use it as a soft guidance for later head-target prediction. To this end, we utilize the head priors encoded in its subsequent head feature , which is derived by passing through a single convolutional layer. After obtaining head priors, is re-injected and fused with via concatenation to provide additional guidance for head localization. The resulting feature, denoted as is then utilized to predict head-target proposals.
3.3 Visual Association via Connection Maps
@@
Given the scene feature for locating potential heads and targets, we propose a novel visual association scheme via multi-task learning as the model not only predicts heads and targets but also connection maps as heatmaps. Taking as input, a prediction head with convolutional layers generates head-target proposals. Each proposal comprises three types of heatmaps: head heatmaps , the corresponding gaze heatmaps , and the connection maps . In contrast to previous approaches that typically link heads and gaze targets through a gaze vector prediction or the predefined field of view parameters [16, 13, 40, 35], the proposed connection maps C explicitly learn the link between the heads and targets. This is motivated by the fact that the prediction of head heatmaps H and gaze heatmaps G are isolated, so it is essential to have these connection maps C to force the model to learn their associations. It is important to note that the connection maps C are a unidirectional mapping that only focus on highlighting the connection between the head and its corresponding target. Through the use of three-heatmap prediction structure, the model is encouraged to learn the individual features of heads and gaze targets, as well as their associations.
3.4 Out-of-frame Flag
@@
Similar to previous gaze target detection works [41, 40, 35], GazeHTA predicts an out-of-frame flag to indicate whether the person is looking outside the frame, in which case the predicted gaze targets are not considered. Same as the head-target proposals, the network utilizes to predict out-of-frame flags, denoted as , using a convolutional layer followed by a single fully-connected layer. Although we can use the feature from a later stage, we empirically found that using for the out-of-frame prediction yields the best overall performance in the final head-target instance prediction. This suggests that out-of-frame prediction may offer early guidance, thereby enhancing the accuracy of head-target proposals. Additionally, this design feature offers practical benefits as it enables the network to skip the computation of head-target proposals during inference if the predicted gaze target is out-of-frame.
3.5 Training Objective
@@
4 Experiments
4.1 Datasets
@@
VideoAttentionTarget [9] is a video-based gaze target dataset. It comprises 71,666 frames from 1,331 clips, with 164,514 annotations on head-target instances. Following the training procedure outlined in [41, 39, 40], we adopt a sampling strategy on the training data, where one frame is selected from every five consecutive frames to avoid overfitting. We sampled 5769 samples from the training set for validation purposes. The test set of this dataset consists of 2,771 images with 6,377 annotations on head-target instances that remain unchanged in our experiments. The data contain human faces and were retrieved from YouTube videos covering various sources including live interviews, sitcoms, reality shows, and movie clips. There was no participant consent form since the data are from public data.
4.2 Ground-truth Generation.
@@ 4.3 Implementation Details
We use the pre-trained image encoder and denoising U-Net from Stable Diffusion [27] as the backbone. During training, the denoising U-Net is finetuned with a learning rate set to 0.1 of the base learning rate, while the image encoder is frozen without updating. We utilize AdamW optimizer [18] along with a one-cycle-policy learning rate scheduler [33]. For the GazeFollow dataset, the maximum learning rate is set to 3.2e-5 with 20 epochs of training. For the VideoAttentionTarget dataset, the maximum learning rate is set to 4e-4 with 10 epochs of training. The weights for training loss , , , , are set as 1.0, 2.5, 1.0, 1.0, and 1.0, respectively. The number of head-target proposals is set to 20 following previous works [39, 41]. The output heatmap width and height are both set to .
4.4 Evaluation Metrics
@@
Following previous works [9, 25, 40, 41], we use the Area Under Curve (AUC) and gaze point distance (Dist.) metrics for evaluation. These two metrics are assessed exclusively when the ground-truth gaze targets are labeled as “in-frame”. AUC indicates the confidence level of the predicted gaze target locations with respect to the ground-truths. Gaze point distance measures the Euclidean distance between the predicted gaze points and the ground-truths, and it is normalized by the image size. The predicted gaze points are inferred from the location with the maximum value in the gaze heatmaps G. For the GazeFollow dataset, both the average distance (Avg. Dist.) and minimum distance (Min. Dist.) are reported due to multiple annotation instances for each given head.
@@
We employ mean Average Precision (mAP) to assess the predicted head-target instances, following the setting in [41]. A head-target instance is regarded as true positive if and only if the Intersection-Over-Union (IOU) ratio between the predicted human head box and ground-truth is greater than 0.5, and the normalized gaze distance is less than 0.15. We perform heatmap thresholding [21] followed by contour finding on H to extract the head bounding boxes. If gaze targets are labeled as out-of-frame, only the head IOU would be considered.
@@
As for the out-of-frame gaze target instances in the VideoAttentionTarget dataset, we report Average Precision (AP) on the out-of-frame flags in our evaluation. This metric does not apply to the GazeFollow dataset since the test set only includes in-frame gaze targets.
4.5 Comparison with State-of-the-Art
Method GazeFollow VideoAttentionTarget AUC Avg. Dist. Min. Dist. mAP AUC Dist. AP mAP VideoAttention [9] 0.921 0.137 0.077 - 0.860 0.134 0.853 - DAM [13] 0.922 0.124 0.067 - 0.905 0.108 0.896 - ESCNet [2] 0.928 0.122 - - 0.885 0.120 0.869 - HGTTR [41] 0.917 0.133 0.069 0.547 0.893 0.137 0.821 0.514 Tonini et al. [40] 0.922 0.072 0.033 0.573 0.923 0.102 0.944 0.607 GazeHTA 0.935 0.062 0.025 0.639 0.951 0.069 0.999 0.762

### v3

@@ Abstract
Precisely detecting which object a person is paying attention to is critical for human-robot interaction since it provides important cues for the next action from the human user. We propose an end-to-end approach for gaze target detection: predicting a head-target connection between individuals and the target image regions they are looking at. Most of the existing methods use independent components such as off-the-shelf head detectors or have problems in establishing associations between heads and gaze targets. In contrast, we investigate an end-to-end multi-person Gaze target detection framework with Heads and Targets Association (GazeHTA), which predicts multiple head-target instances based solely on input scene image. GazeHTA addresses challenges in gaze target detection by (1) leveraging a pre-trained diffusion model to extract scene features for rich semantic understanding, (2) re-injecting a head feature to enhance the head priors for improved head understanding, and (3) learning a connection map as the explicit visual associations between heads and gaze targets. Our extensive experimental results demonstrate that GazeHTA outperforms state-of-the-art gaze target detection methods and two adapted diffusion-based baselines on two standard datasets.
I Introduction
A collaborative robot must quickly and accurately understand the human user intentions [1, 2, 3]. Among non-verbal cues, human eye gaze is a critical indicator of both intention [4] and attention [5]. Eye gaze has been extensively applied in human-robot interaction for tasks such as hand motion prediction [6], robot navigation [7, 8], pick-and-place actions [9, 10], object referencing [11], walking assistance [12], robotic manipulation [13], and facilitating the transfer of objects between humans and robots [14]. In gaze estimation, a common approach is to infer the 3D gaze direction from the eyes [15, 16, 17], face [15, 18, 19], or full-body joints [20]. However, many real-world applications demand not just the gaze direction, but the precise identification of the region or object a person is looking at—referred to as the gaze target.
@@ I Introduction
In contrast to conventional gaze estimation methods, Gaze target detection aims to directly associate individuals and their gaze targets within a single image [21] or across multiple video frames [22], offering an end-to-end solution for human attention estimation. Most of the gaze target detection approaches employ two-stream architectures, where one stream focuses on scene feature extraction while the other one learns head features [23, 24, 25, 26]. These approaches face challenges such as the lack of direct association between heads and gaze targets, and the dependency on off-the-shelf head detectors as demonstrated in recent studies [27, 28]. Moreover, the majority of previous methods are restricted to processing one head at a time, requiring repeated processing to identify gaze targets for all individuals in the scene when multiple people are present. Tu et al. [27] addresses the challenges in the two-stream approaches with an end-to-end multi-person gaze target detection method, to directly predict multiple pairs of head-target instances. Nevertheless, this approach does not establish strong connections between heads and gaze targets, which restricts the final detection accuracy. Tonini et al. [28] improves the association between heads and gaze targets, and achieves state-of-the-art performance. It is, however, not fully end-to-end due to its dependency on preliminary object detection. Pre-training the object detector also limits the object categories considered for gaze targets. This constraint may exclude diverse object classes or regions within the environment that are beyond the scope of the pre-trained object detector.
@@
There are two major challenges in the gaze target detection task: 1) detecting the heads and the gaze targets; and 2) associating heads and gaze targets to predict which head is looking at which target. Addressing the first challenge involves identifying the relevant heads, objects, or other regions as potential gaze targets. This can be solved by using powerful feature extractors, such as diffusion-based models, which have demonstrated remarkable performance in various visual perception tasks [29, 30, 31, 32, 33]. However, effectively integrating these pre-trained large foundation models into gaze target detection remains an open question. The second challenge entails linking heads and corresponding gaze targets. Despite efforts to employ gaze estimation results as constraints on the gaze target locations [24, 34, 28, 35], most existing methods have not fully addressed the association between heads and gaze targets.
@@
In this paper, we propose GazeHTA: a novel end-to-end framework for multi-person gaze target detection. We exploit features from a pre-trained diffusion model, while emphasizing head regions with a novel re-injection method. We then associate the heads and gaze targets explicitly with a connection map in the image. These connection maps serve as links between the heads and gaze targets in the image, offering additional supervision during model training. An overview is shown in Fig. 1. Our experiments demonstrate that GazeHTA achieves state-of-the-art performance while not being restricted to the diffusion model backbone. In summary, the main contributions of this paper include:
@@
Our extensive experiments show that GazeHTA achieves state-of-the-art performances on standard datasets.
II Related Works
II-A Gaze Target Detection
@@
Recently, end-to-end approaches have emerged to simultaneously detect multiple heads and their associated gaze targets [27, 28]. The HGTTR is a transformer-based network that predicts several head-target instances comprising head locations, out-of-frame flags, and gaze heatmaps [27]. Unfortunately, its performance suffers from brittle associations between heads and gaze targets. As a result, the latest advancements leverage the concept of objects to establish associations between heads and gaze targets [28, 38]. These object-aware frameworks initially predict objects and subsequently associate head objects with all detected objects. However, these methods revert to a two-stage approach because of their reliance on detected objects for head-object association. Moreover, the scope of gaze targets is limited to the classes that the pre-trained object detector was trained. To efficiently learn the correlation between heads and gaze targets, it has been shown that explicitly connecting the head and target with a heatmap is feasible [39] while the potential of such an explicit connection has not been explored yet. In contrast to previous approaches, GazeHTA offers an end-to-end solution with implicit target feature extraction, enhanced head priors, and finally, an explicit association between heads and gaze targets through connection maps.
II-B Diffusion Models
Originally from diffusion probabilistic models [40], diffusion models are designed to learn the distribution of data for generating new samples. This process involves iteratively applying transformations from a noise image to a realistic image [31, 41, 42, 43].
@@ II-B Diffusion Models
With the remarkable success of diffusion models in image generation, recent studies showed their potential across various downstream visual perception tasks such as object classification [44] and depth estimation [32], which are strongly correlated with the gaze target detection task. It is straightforward to extend to gaze target detection that the scene image serves as the condition for predicting head and gaze heatmaps through multi-step denoising. However, the output of gaze target detection is heatmaps instead of high-quality images. Therefore, directly applying the denoising process may not be efficient for gaze target detection.
@@
Despite its simplicity, employing a pre-trained diffusion model as a feature extractor is effective for various visual perception tasks, including image segmentation and depth estimation [33, 45]. Considering the output format of gaze target detection is a heatmap which resembles that of image segmentation and depth estimation, namely the spatial maps or heatmaps, they may potentially benefit from the same model architecture. Following the same strategy, GazeHTA employs a pre-trained diffusion model as the backbone to extract both high-level and low-level semantic features for object understanding and head localization, which are crucial for accurate gaze target detection.
III GazeHTA
@@
The pipeline of GazeHTA is shown in Fig. 2. Taking only a scene image as the input, GazeHTA extracts multi-scale scene features for out-of-frame binary predictions and head-target proposals. For the head-target proposal, a head feature is learned and re-injected back to enhance the head priors. Subsequently, heads and gaze targets are predicted in the form of heatmaps. Within each head-target instance, GazeHTA establishes a visual association between the head and its corresponding gaze target by a connection map.
III-A Scene Feature Extraction
@@
In GazeHTA, the backbone to extract scene features is a pre-trained denoising U-Net from Stable Diffusion [31]. The use of Stable Diffusion, a large-scale text-to-image generator, is due to its capability to encompass features with both low-level and high-level semantic knowledge [46] acquired from large-scale real-world data. By taking the scene image as the only input, GazeHTA encodes it with a pre-trained image encoder, followed by a denoising U-Net from the pre-trained diffusion model. We extract intermediate features from four different levels in the U-Net with a similar operation to the VPD [33]. These multi-scale scene features are then encoded into a compact representation for both out-of-frame predictions and head-target proposals. We expect the implicit understanding of the scene features can facilitate the identification of potential heads and gaze targets. Note that, unlike VPD [33], we eliminate the entire text prompt module by removing the cross-attention layers in the denoising U-Net, as it is not necessary nor effective for the gaze target detection task.
III-B Head Feature Re-injection
@@
Accurately locating the human head position is essential for successful gaze target detection. To improve head detection, we introduce supervision on a head detection map, which aims to detect all instances of heads within the scene image. To this end, the multi-scale scene features are passed through a convolution-based decoder that gradually expands the spatial dimension into a decoded feature, denoted as . is then passed through convolutional layers to generate an intermediate head detection map for head prior enhancement. To supervise the training on , we apply the detection results from an off-the-shelf head detector 111https://github.com/SibiAkkash/yolov5-crowdhuman as pseudo labels since the heads are not fully annotated in the training data. Although provides insights into potential head positions, we only use it as a soft guidance for later head-target prediction. To this end, we utilize the head priors encoded in its subsequent head feature , which is derived by passing through a single convolutional layer. After obtaining head priors, is re-injected and fused with via concatenation to provide additional guidance for head localization. The resulting feature, denoted as is utilized to predict head-target proposals.
III-C Visual Association via Connection Maps
@@
Given the scene feature for locating potential heads and targets, we use a visual association scheme via multi-task learning as the model not only predicts heads and targets but also connection maps as heatmaps. Taking as input, a prediction head with convolutional layers generates head-target proposals. Each proposal comprises three types of heatmaps: head heatmaps , the corresponding gaze heatmaps , and the connection maps . In contrast to previous approaches that typically link heads and gaze targets through a gaze vector prediction or the predefined field of view parameters [24, 34, 28, 35], the proposed connection maps C explicitly learn the link between the heads and targets. This is motivated by the fact that the prediction of head heatmaps H and gaze heatmaps G are isolated, so it is essential to have these connection maps C to force the model to learn their associations [39]. It is important to note that the connection maps C are a unidirectional mapping that only focuses on highlighting the connection between the head and its corresponding target.
III-D Out-of-frame Flag
@@
Similar to previous gaze target detection works [27, 28, 35], GazeHTA predicts an out-of-frame flag to indicate whether the person is looking outside the frame, in which case the predicted gaze targets are not considered. Same as the head-target proposals, the network utilizes to predict out-of-frame flags, denoted as , using a convolutional layer followed by a single fully-connected layer. Although we can use the feature from a later stage, we empirically found that using for the out-of-frame prediction yields the best overall performance in the final head-target instance prediction.
III-E Training Objective
@@
IV Experiments
Method | GazeFollow | VideoAttentionTarget | ||||||
AUC | Avg. Dist. | Min. Dist. | mAP | AUC | Dist. | AP | mAP | |
VideoAttention [25] | 0.921 | 0.137 | 0.077 | - | 0.860 | 0.134 | 0.853 | - |
DAM [34] | 0.922 | 0.124 | 0.067 | - | 0.905 | 0.108 | 0.896 | - |
ESCNet [26] | 0.928 | 0.122 | - | - | 0.885 | 0.120 | 0.869 | - |
Sharingan [36] | 0.944 | 0.113 | 0.057 | - | - | 0.107 | 0.891 | - |
HGTTR [27] | 0.917 | 0.133 | 0.069 | 0.547 | 0.893 | 0.137 | 0.821 | 0.514 |
Tonini et al. [28] | 0.922 | 0.072 | 0.033 | 0.573 | 0.923 | 0.102 | 0.944 | 0.607 |
GazeHTA | 0.929 | 0.063 | 0.025 | 0.626 | 0.957 | 0.064 | 1.000 | 0.772 |
IV-A Datasets
@@
VideoAttentionTarget [25] is a video-based gaze target dataset. It comprises 71,666 frames from 1,331 clips, with 164,514 annotations on head-target instances. Following the training procedure outlined in [27, 37, 28], we adopt a sampling strategy on the training data, where one frame is selected from every five consecutive frames to avoid overfitting. We sampled 6 clips from the training set for validation purposes. The test set of this dataset consists of 2,771 frames from 10 clips with 6,377 annotations on head-target instances that remain unchanged in our experiments. The data contain human faces and were retrieved from YouTube videos covering various sources including live interviews, sitcoms, reality shows, and movie clips. There was no participant consent form since the data are from public data.
IV-B Ground-truth Generation.
@@ IV-C Implementation Details
We use the pre-trained image encoder and denoising U-Net from Stable Diffusion [31] as the backbone. During training, the denoising U-Net is finetuned, while the image encoder is frozen without updating. We utilize AdamW optimizer [52] along with a one-cycle-policy learning rate scheduler [53]. For the GazeFollow dataset, the maximum learning rate is set to 3.2e-5 with 20 epochs of training. For the VideoAttentionTarget dataset, the maximum learning rate is set to 4e-4 with 10 epochs of training. The weights for training loss , , , , are set as 1.0, 2.5, 1.0, 1.0, and 1.0, respectively. The number of head-target proposals is set to 20 following previous works [37, 27]. The output heatmap width and height are both set to .
IV-D Evaluation Metrics
@@
Following previous works [25, 21, 28, 27], we use the Area Under Curve (AUC) and gaze point distance (Dist.) metrics for evaluation. These two metrics are assessed exclusively when the ground-truth gaze targets are labeled as “in-frame”. AUC indicates the confidence level of the predicted gaze target locations with respect to the ground-truths. Gaze point distance measures the Euclidean distance between the predicted gaze points and the ground-truths, and it is normalized by the image size. The predicted gaze points are inferred from the location with the maximum value in the gaze heatmaps G.
@@
We employ mean Average Precision (mAP) to assess the predicted head-target instances, following the setting in [27]. A head-target instance is regarded as true positive if and only if the Intersection-Over-Union (IOU) ratio between the predicted and ground-truth head boxes is greater than 0.5, and the normalized gaze distance is less than 0.15. We extract the head bounding box with heatmap thresholding [54] followed by contour finding on H. Only the head IOU is considered if gaze targets are labeled as out-of-frame.
@@
For the GazeFollow dataset, both the average distance (Avg. Dist.) and minimum distance (Min. Dist.) are reported due to multiple annotation instances for each given head. As for the out-of-frame gaze target instances in the VideoAttentionTarget dataset, we report Average Precision (AP) on the out-of-frame flags in our evaluation. This metric does not apply to the GazeFollow dataset since the test set only includes in-frame gaze targets.
IV-E Comparison with State-of-the-Art